OpenStudy (anonymous):

The range of values of x satisfying the inequality \[\left| 2x +3 \right|\ge7 \] is?

OpenStudy (anonymous):

@satellite73 help

OpenStudy (saifoo.khan):

Can we help too? :S

OpenStudy (anonymous):

\[2x+3\geq 7\] or \[2x+3\leq -7\] solve both inequalities for \(x\) you will get two separate intervals saifoo claus can finish

OpenStudy (anonymous):

2x+3≥7

OpenStudy (saifoo.khan):

Haha @satellite73 . @shameer1 there will be two different values.

OpenStudy (anonymous):

i got it as -5 and 2 but the answer is (-infinity,-5) U {2, infinity)

OpenStudy (saifoo.khan):

Try graphing the answers.. then you can understand this.

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

you are solving INEQUALITIES

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no equalities. so your solutions are intervals, not numbers. in other words when you solve \[2x+3\leq -7\] subtract 3 get \[2x\leq -10\] divide by 2 get \[x\leq 5\] that is an inequality, means all numbers less than or equal to -5. as an interval it looks like \[(-\infty,-5)\]

OpenStudy (anonymous):

similarly \[2x+3\geq 7\] \[2x\geq4\] \[x\geq 2\] all numbers greater than or equal to 2 as in interval looks like \[[2,\infty)\]

OpenStudy (anonymous):

actually my first interval was wrong, should have been written as \[(-\infty,-5]\]

OpenStudy (anonymous):

keep the inequality when you solve, do not change it to an equal sign or you will get confused at the end

OpenStudy (anonymous):

hey how that ign changes ist equation